# UniForm: A Unified Multi-Task Diffusion Transformer for Audio-Video Generation

### v3

@@ Abstract.
With the rise of diffusion models, audio-video generation has been revolutionized. However, most existing methods rely on separate modules for each modality, with limited exploration of unified generative architectures. In addition, many are confined to a single task and small-scale datasets. To address these limitations, we first propose UniForm, a unified multi-task diffusion transformer that jointly generates audio and visual modalities in a shared latent space. A single diffusion process models both audio and video, capturing the inherent correlations between sound and vision. Second, we introduce task-specific noise schemes and task tokens, enabling a single model to support multiple tasks, including text-to-audio-video, audio-to-video, and video-to-audio generation. Furthermore, by leveraging large language models and a large-scale text-audio-video combined dataset, UniForm achieves greater generative diversity than prior approaches. Extensive experiments show that UniForm achieves the state-of-the-art performance across audio-video generation tasks, producing content that is both well-aligned and close to real-world data distributions. Our demos are available at https://uniform-t2av.github.io/.
Multimodal Generation, Multi-task Learning, Audio Generation, Video Generation, Diffusion Transformer
ccs: Information systems Multimedia information systemsccs: Computing methodologies Computer visionccs: Computing methodologies Multi-task learningccs: Computing methodologies Natural language processing
Figure 1. Illustration of multimodal-conditioned audio-video generation. Text can create audio-video directly; audio or video can serve as a condition to guide the generation of the other.
[…]
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
1. Introduction
@@
With the flourishing of deep learning, artificial intelligence generated content (AIGC) has revolutionized multimodal creation, enabling vivid generation across text (Chung et al., 2024; Li et al., 2024), images (Rombach et al., 2022; Zhu et al., 2023), audio (Liu et al., 2024b; Wang et al., 2024a), and video (Song et al., 2024; Zheng et al., 2024). This progress has driven innovation in creative industries and expanded the scope of digital media. However, most AIGC systems remain confined to single modality. For example, text-to-video methods, like (Singer et al., 2023) decompose temporal U-Nets into spatial and temporal blocks, (Blattmann et al., 2023) extends latent diffusion to video by modeling time, and (Wang et al., 2024b) uses cascaded diffusion with joint image-video tuning. Despite strong visual results, these methods lack sound, overlooking multisensory integration, which is a key element for immersive experiences.
@@
Recent efforts have begun exploring audio-video co-generation. MM-Diffusion (Ruan et al., 2023) employs dual U-Net subnets for parallel audio-video synthesis. In contrast, another emerging diffusion backbone is the Diffusion Transformer (DiT) (Peebles and Xie, 2023), which has demonstrated remarkable performance in various content generation tasks. Building on this, (Hayakawa et al., 2025) employs two diffusion processes, followed by a joint discriminator to integrate audio and video. Meanwhile, AV-DiT (Wang et al., 2024c) adopts a shared DiT backbone pre-trained exclusively on image data, facilitating audio and video generation through the addition of lightweight adapters.
@@
Although these co-generation methods have shown strong results, there remains room for further exploration. First, they rely on two separate sub-networks to generate audio and video independently, which may limit the depth of cross-modal integration. Second, these methods rely solely on text labels and are mostly trained on small-scale datasets, which limits the diversity of the generated results. Finally, most methods focus solely on the text-to-audio-video (T2AV) task and do not support audio-to-video (A2V) or video-to-audio (V2A) generation. While (Xing et al., 2024) enables the three tasks, each task requires distinct pre-trained models in their work.
@@
Inspired by the natural coupling of sound and vision in real-world videos, we ask whether a unified model can improve alignment and consistency across modalities. In this work, we present the unified multi-task diffusion transformer (UniForm), the unified system for generating audio-video content. Moreover, as shown in Figure 1, all three tasks are handled by a single unified model. Our contributions are summarized as follows:
  • We propose a unified diffusion framework for the synchronous generation of both audio and video content. UniForm concatenates vision and audio tokens, enabling joint diffusion modeling in a unified latent space and implicitly learning cross-modal correlations.
@@
  • We propose a unified backbone for multi-task audio-video generation, including T2AV, A2V, and V2A. We incorporate task-specific noise schemes and task tokens to specify the target task. For the latter two tasks, text prompts can also be optionally used as auxiliary input to enable fine-grained control and enhance performance.
@@
  • We employ a large language model (LLM) to encode the text, a process that does not rely on the text labels used by previous methods, thereby providing finer control for the model and enhancing the diversity of generated content. To fully leverage this advantage , we produced an extensive caption corpus to facilitate model training.
@@
  • Experiments show that UniForm achieves performance comparable to the state-of-the-art single-task baselines, both in objective metrics and in subjective audio-visual quality. Remarkably, this performance is achieved without fine-tuning on task-specific datasets, as the model is trained solely in a multi-task setting. In addition, compared to non-unified methods (i.e., using separate backbones for each task (Xing et al., 2024)), our approach demonstrates consistent advantages across the board.
2. Related Work
2.1. Video to Audio Generation
@@
In this paper, we focus on “Foley” audio, [^1] YouTube: The Magic of Making Sound which refers to sound effects added during post-production to enrich the auditory experience of multimedia (Choi et al., 2023), such as the crunch of leaves or the clinking of glass bottles. Earlier AI-based Foley generation methods were conditioned on class labels (Liu et al., 2021) or text prompts (Liu et al., 2023). Building on this, recent work has expanded video-to-audio generation. SpecVQGAN (Iashin and Rahtu, 2021) adopts a transformer that generates high-fidelity spectrograms from video frames using a VQGAN-based codebook and perceptual audio loss. (Du et al., 2023) requires both silent video and conditional audio to produce candidate tracks, which are filtered using an audio-visual synchronization model. Diff-Foley (Luo et al., 2024) only requires silent video as input. It first aligns audio-visual features through contrastive pretraining (CAVP), then trains a diffusion model on spectrogram latents conditioned on CAVP features. FoleyCrafter (Zhang et al., 2024) introduces optional text prompts for finer control and incorporates a semantic adapter and temporal controller for improved alignment. V-AURA (Viertola et al., 2025) proposes an autoregressive model with high-frame-rate visual encoding and cross-modal fusion for fine-grained temporal precision. VATT (Liu et al., 2024a) presents a multi-modal system that generates audio from video with optional text prompts. Departing from GANs and diffusion, FRIEREN (Wang et al., 2024d) uses flow matching as its generative backbone. To better preserve temporal structure, it introduces a non-autoregressive vector field estimator without temporal downsampling.
2.2. Audio to Video Generation
@@
Given the high information density of video, video-to-audio generation tasks typically treat video as the primary input, with text serving as an auxiliary cue. In contrast, audio to video generation mainly relies on audio alignment. Due to the limited semantic context provided by audio itself, it is difficult for both humans and machines to distinguish (e.g., distinguishing a given audio clip corresponds to climbing stairs or tap dancing). As a result, audio-to-video generation often depends on text or images to supply the missing context, and dedicated efforts in this direction remain relatively limited. (Lee et al., 2022) learns to generate semantically aligned video from audio. It maps audio into the StyleGAN latent space and refines the output using CLIP-based multimodal embeddings to reinforce audio-visual coherence. TPoS (Jeong et al., 2023) follows a stage-wise strategy: it first generates an initial frame from a text prompt, then progressively adapts the visuals based on the audio input. (Yariv et al., 2024) proposes a lightweight adaptor that translates audio features into the input format expected by a pre-trained text-to-video model, enabling audio-driven video generation with minimal changes to the backbone.
@@
Figure 2. Overview of the proposed UniForm. Vision tokens and audio tokens are integrated and processed within a unified latent space using a DiT model to learn their representations. During training, one of three tasks is randomly selected in each iteration, with task tokens guiding the learning of the DiT. The text encoder, the encoder-decoder for video and audio, and the audio vocoder are all pre-trained models that remain frozen throughout.
2.3. Diffusion-based Generation
@@
Diffusion models, as a class of probabilistic generative models, have received growing attention for their remarkable performance in diverse domains such as image generation (Ramesh et al., 2022) and audio synthesis (Liu et al., 2024b). The majority of existing approaches are built upon Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020), which form the foundation of this paradigm. The key idea of diffusion modeling is to define a forward process that progressively perturbs data into Gaussian noise through a sequence of noise-adding steps. The model is then trained to approximate the reverse process, which starts from pure noise and performs iterative denoising steps to recover samples that approximate the original data distribution. To reduce the computational cost of operating in high-dimensional spaces, Latent Diffusion Models (LDMs) (Rombach et al., 2022) shift the diffusion process into a lower-dimensional latent space, enabling more efficient generation. (Peebles and Xie, 2023) explore replacing the previously used U-Net backbone with a transformer operating in latent space. Their results show that, given sufficient compute, the Diffusion Transformer (DiT) produces samples that closer to the original data distribution. In this work, we adopt DiT as the backbone of our multi-task architecture, leveraging its scalability and strong performance in modeling a unified latent space across audio and visual modalities.
3. Method
@@
In this section, we first define the three types of audio-video generation tasks addressed in Section 3.1. Next, we review the preliminary knowledge of diffusion-based generation in Section 3.2. Finally, we present a detailed introduction of the proposed UniForm in Section 3.3.
3.1. Problem Definition
@@
(1) $\hat{z}^{a}_{0},\hat{z}^{v}_{0}=f(z^{a}_{T},z^{v}_{T},c),$
Based on Eq. (1), we introduce three sets of task-specific noise schemes. In the T2AV task, all three inputs are provided. In the A2V task, the audio noise input is removed by setting $z^{a}_{T}=0$; similarly, in the V2A task, we set $z^{v}_{T}=0$. During training, we integrate a classifier-free guidance (CFG) strategy (Ho and Salimans, 2021), conditioning on the text embedding $c$ with a 50% probability. This allows the model to learn various combinations of modality-specific priors within a single unified framework for multimodal generation.
3.2. Diffusion Model
@@
(2) $q(z^{k}_{t}|z^{k}_{t-1})=\mathcal{N}\left(\sqrt{1-\beta_{t}}\,z^{k}_{t-1},\beta_{t}\mathbf{I}\right),k\in\{a,v\},$
(3) $q(z^{k}_{t}|z^{k}_{0})=\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\,z^{k}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right),k\in\{a,v\},$
where $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. A reparameterization method (Song et al., 2021) simplifies the sampling of any intermediate states $z^{a}_{n}$ and $z^{v}_{n}$ from the initial states $z^{a}_{0}$ and $z^{v}_{0}$, using the following formulation:
(4) $z^{k}_{t}=\sqrt{\bar{\alpha}_{t}}\,z^{k}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon^{k},k\in\{a,v\},$
@@
(5) $\mathcal{L}=\gamma_{t}\sum_{k\in\{a,v\}}\mathbb{E}_{\epsilon^{k}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),z^{k}_{0}}\left\|\epsilon^{k}_{t}-\epsilon_{\theta}^{(t)}(z^{k}_{t},c)\right\|_{2}^{2},$
@@
(6) $p_{\theta}(z^{k}_{0:T}|c)=p(z^{k}_{T})\prod_{t=1}^{T}p_{\theta}(z^{k}_{t-1}|z^{k}_{t},c),$
(7) $p_{\theta}(z^{k}_{t-1}|z^{k}_{t},c)=\mathcal{N}\left(\mu_{\theta}^{(t)}(z^{k}_{t},c),\tilde{\beta}^{(t)}\right),$
(8) $\mu_{\theta}^{(t)}(z^{k}_{t},c)=\frac{1}{\sqrt{\alpha_{t}}}\left[z^{k}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}^{(t)}(z^{k}_{t},c)\right],$
(9) $\tilde{\beta}^{(t)}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t},$
@@
3.3. UniForm
3.3.1. Video & Audio Latent Encoding
As mentioned earlier, we adopt DiT to model a unified latent space shared across audio and visual modalities. FLAN-T5 (Raffel et al., 2020; Chung et al., 2024) is used as the text encoder, where T5 is a high-capacity pretrained LLM known for its strong semantic understanding. During training, the video input is represented as $V\in\mathbb{R}^{T^{v}\times C^{v}\times H\times W}$, where $T^{v}$ denotes the number of temporal frames, each with $C^{v}$ channels, height $H$, and width $W$. A pre-trained video encoder (Zheng et al., 2024) is adopted to extract vision latent $z^{v}_{0}\in\mathbb{R}^{\hat{C}^{v}\times\hat{T}^{v}\times\hat{H}\times\hat{W}}$ from videos, where $\hat{C}^{v}$, $\hat{T}^{v}$, $\hat{H}$, $\hat{W}$ are hidden dimensions of vision tokens. For the audio input, we first apply the Short-Time Fourier Transform (STFT) to convert the waveform from the time domain to the frequency domain. Then, a set of Mel-scale filters is used to generate the Mel spectrogram with shape $\mathbb{R}^{T^{a}\times F}$, where $T^{a}$ denotes the number of temporal frames and $F$ is the number of frequency bins. These Mel spectrograms are subsequently passed through a pre-trained audio VAE encoder (Liu et al., 2023) to obtain the audio latent tokens $z^{a}_{0}\in\mathbb{R}^{\hat{C}^{a}\times\hat{T}^{a}\times\hat{F}\times 1}$. Latent tokens from both modalities first undergo reshaping operations to align their dimensions. Subsequently, these adjusted tokens are concatenated along the last dimension, forming a unified representation that serves as the input to the shared DiT.
3.3.2. Multitask Modeling

### v5

@@ Abstract
With the rise of diffusion models, audio-video generation has been revolutionized. However, most existing methods rely on separate modules for each modality, with limited exploration of unified generative architectures. In addition, many are confined to a single task and small-scale datasets. To overcome these limitations, we introduce UniForm, a unified multi-task diffusion transformer that generates both audio and visual modalities in a shared latent space. By using a unified denoising network, UniForm captures the inherent correlations between sound and vision. Additionally, we propose task-specific noise schemes and task tokens, enabling the model to support multiple tasks with a single set of parameters, including video-to-audio, audio-to-video and text-to-audio-video generation. Furthermore, by leveraging large language models and a large-scale text-audio-video combined dataset, UniForm achieves greater generative diversity than prior approaches. Experiments show that UniForm achieves performance close to the state-of-the-art single-task models across three generation tasks, with generated content that is not only highly aligned with real-world data distributions but also enables more diverse and fine-grained generation.
Index Terms:
Text-to-spatial-audio, audio generation, latent diffusion model.
I Introduction
@@
With the flourishing of deep learning, artificial intelligence generated content (AIGC) has revolutionized multimodal creation, enabling vivid generation across text [1, 2], images [3, 4], audio [5, 6], and video [7, 8]. This progress has driven innovation in creative industries and expanded the scope of digital media. However, most AIGC systems remain confined to single modality. For example, text-to-video methods, like [9] decompose temporal U-Nets into spatial and temporal blocks, [10] extends latent diffusion to video by modeling time, and [11] uses cascaded diffusion with joint image-video tuning. Despite strong visual results, these methods lack sound, overlooking multisensory integration, which is a key element for immersive experiences.
@@
Recent efforts have begun exploring audio-video co-generation. MM-Diffusion [12] employs dual U-Net subnets for parallel audio-video synthesis. Subsequently, MM-LDM [13] utilizes two separate diffusion model to independently process audio and video, ultimately enabling multimodal interaction within a shared high-level semantic space. In contrast, another emerging diffusion backbone is the Diffusion Transformer (DiT) [14], which has demonstrated remarkable performance in various content generation tasks. Building on this, [15] employs two diffusion processes, followed by a joint discriminator to integrate audio and video. Meanwhile, AV-DiT [16] adopts a shared DiT backbone pre-trained exclusively on image data, facilitating audio and video generation through the addition of lightweight adapters.
Figure 1: Illustration of multimodal-conditioned audio-video generation. Text can create audio-video directly; audio or video can serve as a condition to guide the generation of the other.
@@
Although these co-generation methods have shown strong results, there remains room for further exploration. First, they rely on two separate sub-networks to generate audio and video independently, which may limit the depth of cross-modal integration. Second, these methods rely solely on text labels and are mostly trained on small-scale datasets, which limits the diversity of the generated results. Finally, most methods focus solely on the text-to-audio-video (T2AV) task and do not provide adequate exploration and experimentation on other related generation tasks, such as audio-to-video (A2V) or video-to-audio (V2A) generation. While [17] enables the three tasks, each task requires distinct pre-trained models in their work.
@@
Inspired by the natural coupling of sound and vision in real-world videos, we ask whether a unified model can improve alignment and consistency across modalities. In this work, we propose Uniform, a unified multi-task diffusion transformer for generating consistent audio-video content. Moreover, as shown in Figure 1, all three tasks are handled by a unified network with shared weights. Our contributions are summarized as follows:
  • We propose a unified DiT to jointly generate audio and video content synchronously. UniForm constructs a unified latent representation space by concatenating the latent features encoded from the audio Variational Autoencoder (VAE) and the video VAE, enabling joint diffusion modeling across modalities and implicitly capturing the correlations between modalities.
@@
  • We propose a unified denoising network that can perform multi-task audio-video generation with a single set of model parameters, including V2A, A2V and T2AV. We incorporate task-specific noise schemes and task tokens to specify the target task. For the latter two tasks, text prompts can also be optionally used as auxiliary input to enable fine-grained control and enhance performance.
@@
  • Our method enables the generation of more diverse audio and video content. We use a large language model (LLM) to encode the text, a process that does not rely on the text labels used in previous methods, thereby providing finer control for the model and enhancing the diversity of the generated content. To fully leverage this advantage , we produced an extensive caption corpus to facilitate model training.
@@
  • Experiments show that UniForm achieves performance comparable to the state-of-the-art single-task baselines. Remarkably, this performance is achieved without fine-tuning on task-specific datasets, as the model is trained solely in a multi-task setting. In addition, compared to non-unified methods (i.e., using separate backbones for each task [17]), our approach demonstrates consistent advantages across the board.
The rest of this paper is organized as follows. Section II discusses related work. Section III describes the proposed method in detail. Section IV presents the experimental setup and results. Finally, Section V concludes the study.
II Related Work
II-A Video to Audio Generation
@@
In this paper, we focus on “Foley” audio, [^1] YouTube: The Magic of Making Sound which refers to sound effects added during post-production to enrich the auditory experience of multimedia [18], such as the crunch of leaves or the clinking of glass bottles. Earlier AI-based Foley generation methods were conditioned on class labels [19] or text prompts [20]. Building on this, recent work has expanded video-to-audio generation. SpecVQGAN [21] adopts a transformer that generates high-fidelity spectrograms from video frames using a VQGAN-based codebook and perceptual audio loss. [22] requires both silent video and conditional audio to produce candidate tracks, which are filtered using an audio-visual synchronization model. Diff-Foley [23] only requires silent video as input. It first aligns audio-visual features through contrastive pretraining (CAVP), then trains a diffusion model on spectrogram latents conditioned on CAVP features. FoleyCrafter [24] introduces optional text prompts for finer control and incorporates a semantic adapter and temporal controller for improved alignment. V-AURA [25] proposes an autoregressive model with high-frame-rate visual encoding and cross-modal fusion for fine-grained temporal precision. VATT [26] presents a multi-modal system that generates audio from video with optional text prompts. Departing from GANs and diffusion, FRIEREN [27] uses flow matching as its generative backbone. To better preserve temporal structure, it introduces a non-autoregressive vector field estimator without temporal downsampling.
II-B Audio to Video Generation
@@
Given the high information density of video, video-to-audio generation tasks typically treat video as the primary input, with text serving as an auxiliary cue. In contrast, audio to video generation mainly relies on audio alignment. Due to the limited semantic context provided by audio itself, it is difficult for both humans and machines to distinguish (e.g., distinguishing a given audio clip corresponds to climbing stairs or tap dancing). As a result, audio-to-video generation often depends on text or images to supply the missing context, and dedicated efforts in this direction remain relatively limited. [28] learns to generate semantically aligned video from audio. It maps audio into the StyleGAN latent space and refines the output using CLIP-based multimodal embeddings to reinforce audio-visual coherence. TPoS [29] follows a stage-wise strategy: it first generates an initial frame from a text prompt, then progressively adapts the visuals based on the audio input. [30] proposes a lightweight adaptor that translates audio features into the input format expected by a pre-trained text-to-video model, enabling audio-driven video generation with minimal changes to the backbone.
@@
Figure 2: Overview of the proposed UniForm. Vision tokens and audio tokens are integrated and processed within a unified latent space using a DiT model to learn their representations. During training, one of three tasks is randomly selected in each iteration, with task tokens guiding the learning of the DiT. The text encoder, the encoder-decoder for video and audio, and the audio vocoder are all pre-trained models that remain frozen throughout.
II-C Diffusion-based Generation
@@
Diffusion models, as a class of probabilistic generative models, have received growing attention for their remarkable performance in diverse domains such as image generation [31] and audio synthesis [5]. The majority of existing approaches are built upon Denoising Diffusion Probabilistic Models (DDPMs) [32], which form the foundation of this paradigm. The key idea of diffusion modeling is to define a forward process that progressively perturbs data into Gaussian noise through a sequence of noise-adding steps. The model is then trained to approximate the reverse process, which starts from pure noise and performs iterative denoising steps to recover samples that approximate the original data distribution. To reduce the computational cost of operating in high-dimensional spaces, Latent Diffusion Models (LDMs) [3] shift the diffusion process into a lower-dimensional latent space, enabling more efficient generation. [14] explore replacing the previously used U-Net backbone with a transformer operating in latent space. Their results show that, given sufficient compute, the Diffusion Transformer (DiT) produces samples that closer to the original data distribution. In this work, we adopt DiT as the backbone of our multi-task architecture, leveraging its scalability and strong performance in modeling a unified latent space across audio and visual modalities.
III Method
@@
In this section, we first define the three types of audio-video generation tasks addressed in Section III-A. Next, we review the preliminary knowledge of diffusion-based generation in Section III-B. Finally, we present a detailed introduction of the proposed UniForm in Section III-C.
III-A Problem Definition
@@
$\hat{z}^{a}_{0},\hat{z}^{v}_{0}=f(z^{a}_{T},z^{v}_{T},c),$ (1)
Based on Eq. (1), we introduce three sets of task-specific noise schemes. In the T2AV task, all three inputs are provided. In the A2V task, the audio noise input is removed by setting $z^{a}_{T}=0$; similarly, in the V2A task, we set $z^{v}_{T}=0$. During training, we integrate a classifier-free guidance (CFG) strategy [33], conditioning on the text embedding $c$ with a 50% probability. This allows the model to learn various combinations of modality-specific priors within a single unified framework for multimodal generation.
III-B Diffusion Model
@@
$q(z^{k}_{t}|z^{k}_{t-1})=\mathcal{N}\left(\sqrt{1-\beta_{t}}\,z^{k}_{t-1},\beta_{t}\mathbf{I}\right),k\in\{a,v\},$ (2)
$q(z^{k}_{t}|z^{k}_{0})=\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\,z^{k}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right),k\in\{a,v\},$ (3)
where $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. A reparameterization method [34] simplifies the sampling of any intermediate states $z^{a}_{n}$ and $z^{v}_{n}$ from the initial states $z^{a}_{0}$ and $z^{v}_{0}$, using the following formulation:
$z^{k}_{t}=\sqrt{\bar{\alpha}_{t}}\,z^{k}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon^{k},k\in\{a,v\},$ (4)
@@
$\mathcal{L}=\gamma_{t}\sum_{k\in\{a,v\}}\mathbb{E}_{\epsilon^{k}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),z^{k}_{0}}\left\|\epsilon^{k}_{t}-\epsilon_{\theta}^{(t)}(z^{k}_{t},c)\right\|_{2}^{2},$ (5)
@@
$p_{\theta}(z^{k}_{0:T}|c)=p(z^{k}_{T})\prod_{t=1}^{T}p_{\theta}(z^{k}_{t-1}|z^{k}_{t},c),$ (6)
$p_{\theta}(z^{k}_{t-1}|z^{k}_{t},c)=\mathcal{N}\left(\mu_{\theta}^{(t)}(z^{k}_{t},c),\tilde{\beta}^{(t)}\right),$ (7)
$\mu_{\theta}^{(t)}(z^{k}_{t},c)=\frac{1}{\sqrt{\alpha_{t}}}\left[z^{k}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}^{(t)}(z^{k}_{t},c)\right],$ (8)
$\tilde{\beta}^{(t)}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t},$ (9)
@@
III-C UniForm
III-C1 Video & Audio Latent Encoding
As mentioned earlier, we adopt DiT to model a unified latent space shared across audio and visual modalities. FLAN-T5 [35, 1] is used as the text encoder, where T5 is a high-capacity pretrained LLM known for its strong semantic understanding. During training, the video input is represented as $V\in\mathbb{R}^{T^{v}\times C^{v}\times H\times W}$, where $T^{v}$ denotes the number of temporal frames, each with $C^{v}$ channels, height $H$, and width $W$. A pre-trained video encoder [8] is adopted to extract vision latent $z^{v}_{0}\in\mathbb{R}^{\hat{C}^{v}\times\hat{T}^{v}\times\hat{H}\times\hat{W}}$ from videos, where $\hat{C}^{v}$, $\hat{T}^{v}$, $\hat{H}$, $\hat{W}$ are hidden dimensions of vision tokens. For the audio input, we first apply the Short-Time Fourier Transform (STFT) to convert the waveform from the time domain to the frequency domain. Then, a set of Mel-scale filters is used to generate the Mel spectrogram with shape $\mathbb{R}^{T^{a}\times F}$, where $T^{a}$ denotes the number of temporal frames and $F$ is the number of frequency bins. These Mel spectrograms are subsequently passed through a pre-trained audio VAE encoder [20] to obtain the audio latent tokens $z^{a}_{0}\in\mathbb{R}^{\hat{C}^{a}\times\hat{T}^{a}\times\hat{F}\times 1}$. Latent tokens from both modalities first undergo reshaping operations to align their dimensions. Subsequently, these adjusted tokens are concatenated along the last dimension, forming a unified representation that serves as the input to the shared DiT.
III-C2 Multitask Modeling